OpenStudy (shesolitt):

I WILL MEDAL AND FAN Rosa is saving money for a trip. She has already saved $250 and plans to save $40 per week. This situation can be represented by the function f(x) = 250 + 40x, where x is the number of weeks. How much money has Rosa saved after 6 weeks? $490 $480 $250 $240

OpenStudy (shesolitt):

@jackthegreatest

563blackghost (563blackghost):

Plug in 6 as `x`..... \(\huge\bf{f(x)=250 + 40(6)}\)

OpenStudy (jackthegreatest):

@563blackghost is right

OpenStudy (jackthegreatest):

let me know your answer

OpenStudy (shesolitt):

240?

OpenStudy (jackthegreatest):

250 + 240 =?

563blackghost (563blackghost):

Add the two as jackthegreatest said.

OpenStudy (pinkgirl56):

250 + 40(6) = 250 + 240 = 490 a. 490

OpenStudy (jackthegreatest):

yep correct

OpenStudy (jackthegreatest):

pinkgirl got it

OpenStudy (shesolitt):

thanks eveyonee